# Understanding Token Probability Encoding in Output Embeddings

### v1

@@ Abstract
In this paper, we investigate the output token probability information in the output embedding of language models. We provide an approximate common log-linear encoding of output token probabilities within the output embedding vectors and demonstrate that it is accurate and sparse when the output space is large and output logits are concentrated. Based on such findings, we edit the encoding in output embedding to modify the output probability distribution accurately. Moreover, the sparsity we find in output probability encoding suggests that a large number of dimensions in the output embedding do not contribute to causal language modeling. Therefore, we attempt to delete the output-unrelated dimensions and find more than 30% of the dimensions can be deleted without significant movement in output distribution and degeneration on sequence generation. Additionally, in training dynamics, we use such encoding as a probe and find that the output embeddings capture token frequency information in early steps, even before an obvious convergence starts.
@@
Hakaze Cho1,∗, Yoshihiro Sakai1, Kenshiro Tanaka1, Mariko Kato1, Naoya Inoue1,2,∗ 1Japan Advanced Institute of Science and Technology, 2RIKEN {yfzhao, naoya-i}@jaist.ac.jp
@@ 1 Introduction
Modern Language Models (LMs) have two kinds of embeddings. One is the input embedding $E^{(i)}$ located at the earliest layer of LMs, for mapping the input token index into distributed inner representation. The other is the output embedding $E^{(o)}$ in the Language Modeling head (LM head), for mapping the hidden state to the predicted probability distribution of the next token in the causal language modeling task.
@@
Since the output embeddings were often tied with the input embeddings Chung et al. (2020); Press and Wolf (2017), i.e. input embeddings are directly used as the output embedding, the behaviors and features of independent output embeddings are rarely investigated. Along with the scaling of LMs, such embedding tying, which is proven to be harmful to model performance Chung et al. (2020), is gradually being unused in modern LMs such as GPT-J Wang and Komatsuzaki (2021) and LLaMa 2 Touvron et al. (2023). This raises attention to the output embeddings, and explaining its underlying mechanism can be beneficial to understanding and improving LMs.
@@
Figure 1: The PCA result of the output embedding parameters of GPT2. Colors refer to the ranking percentile of the averaged output token probability. Output embeddings encode the probabilities linear-likely.
@@
The most obvious and expected role of the LM heads is to map the last hidden states into token probabilities. Therefore, following Kobayashi et al. (2023), who found an encoding of the averaged output token probability distribution in the bias term of the output LM head, this paper also focuses on the averaged output probabilities [^1] We may omit the ”averaged” in the following. as an overall representation of LM outputs. We observe there is a linear-like correlation between the output token probabilities and the output embedding as shown in Fig. 1. In §2, our mathematical derivation indicates that $\mathrm{softmax}$ LM head naturally encodes the output probabilities log-linearly in a common direction of the output embeddings, as long as the output dimension is sufficiently large, and the output values are concentrated.
@@
To prove our derivation empirically, we conduct Multiple Linear Regression (MLR) on output probabilities against output embeddings, where a strong log-linear correlation is observed. Additionally, we find that: (1). Almost all directions highly correlated with output probability are the top principal components of the embedding matrix; (2). Only a few dimensions of output embedding are related to output probability.
@@
Based on such findings, to further demonstrate the possible applications, in §3, we try to edit the output probability by the output embedding. We use a linear and local vectorized algorithm, modifying a small portion of dimensions along the encoding direction in the output embeddings for editing the probabilities. Our experiments find that: even on embedding-tied models, our editing method has respectable precision and a large usable range, stably scaling the probability of tokens up to 20x (both scaling up or down), with little disturbing on the normal prediction process of LMs. Moreover, with the encoding direction estimated on few-shot examples, the editing remains precise. Such results suggest that such a log-linear correlation we found is with good stability and generalization.
@@
Moreover, the aforementioned phenomenon demonstrates that most of the dimensions of the output embedding have minor effects on output probability. Therefore, we try removing these dimensions to reduce the parameters in output embedding without obvious harm to the causal language modeling in §4. Our experiments show that more than 30% of the output embedding dimensions can be removed without significant impairment.
(a) GPT2, 137M
(b) GPT-J, 6B
Figure 2: The MLR results on GPT2 and GPT-J.
@@
The contribution of this paper can be summarized as:
  • We find a log-linear correlation as an encoding of the output probability in the output embedding. That is, the output token probabilities are encoded in a particular common direction on the output embeddings log-linearly.
  • Based on the findings, we edit the averaged output probabilities using such an encoding.
  • Based on the findings, we remove dimensions with weak correlation to output probabilities without harm to the LMs.
  • Based on the findings, we find that the LMs learn the token frequency in the training corpus at very early training steps.
@@
Model GPT2 GPT2-XL Pythia GPT-J
Parameter # 137M 1.6B 2.8B 6B
Embedding tied -- --
Random Adj.$R^{2}$ $0.001$ $0.000$ $0.001$ $0.000$
Adj.$R^{2}$ on $E^{(o)}_{w}$ $0.892$ $0.893$ $0.856$ $0.882$
Adj.$R^{2}$ on $E^{(i)}_{w}$ $(0.892)$ $(0.893)$ $0.814$ $0.658$
(a) GPT2, 137M, PCA
(b) GPT2-XL, 1.6B, PCA
(c) Pythia, 2.8B, PCA
(d) GPT-J, 6B, PCA
@@
Figure 3: Only a few directions/dimensions of output embedding are strongly correlated to the output probabilities. (a-d): horizontal axis: the principle components of output embedding, vertical axis: absolute Spearman $r$ between the principle and the output probability distribution, color bar: the variance ratio loaded in the principle component; (e-h): horizontal axis: original dimensions, vertical axis: absolute MLR slopes between this dimension and the output probability distribution, color bar: the absolute Spearman correlations on the dimension.
@@ 2 Token Probability Encoding in Output Embedding
In this section, we preliminarily reveal how the output probabilities are encoded in the output embedding by derivation, then conduct simple numerical experiments to confirm it empirically.
Input: Language model $P_{\theta}(x)$ with output embedding $E^{(o)}$; Probability detecting dataset $\mathcal{D}$;
Token index to be edited $w$; Expected scale of the edited token’s probability $r$;
Editing amount allocation parameter $b$
@@
$\mathbf{\alpha}_{\mathcal{D},\theta}\leftarrow|\mathcal{D}|^{-1}\sum_{x\in\mathcal{D}}P_{\theta}(x)$ ;
  // Calculate the averaged probability distribution
@@
$A_{\mathcal{D}},\mathcal{S}\leftarrow\mathrm{MLR}(-\log\mathbf{\alpha}_{\mathcal{D},\theta},E^{(o)})$ ;
  // Conduct the MLR, get the slope and confidence
@@
$\Omega\leftarrow|A_{\mathcal{D}}|^{b}\times\mathcal{S}\times\||A_{\mathcal{D}}|^{b}\times\mathcal{S}\|_{1}^{-1}$ ;
  // Allocate the editing weight of dimensions
@@
  // Apply the editing to output embedding
@@
10return Updated LM $P_{\theta^{\prime}}(x)$ with new output embedding $E_{w}^{{}^{\prime}(o)}$
2 $\mathbf{\alpha}_{\mathcal{D},\theta}\in\mathbb{R}^{|\mathbb{V}|}$: the output token probability distribution among the vocabulary; $|\mathcal{D}|$: the length of the probability detecting dataset; $\mathcal{S}$: the element-wise confidence ($p$-value) of the MLR;
@@
Algorithm 1 Output token probability editing.
4 $\mathbf{\alpha}_{\mathcal{D},\theta}\in\mathbb{R}^{|\mathbb{V}|}$: the output token probability distribution among the vocabulary; $|\mathcal{D}|$: the length of the probability detecting dataset; $\mathcal{S}$: the element-wise confidence ($p$-value) of the MLR;
@@ 2.1 Mathematical Log-linear Form
$P_{\theta}(w|x)=\frac{e^{E_{w}^{(o)}\cdot h_{x}}}{\sum_{i\in\mathbb{V}}e^{E_{i}^{(o)}\cdot h_{x}}},$
@@
$-\log[P_{\theta}(w|x)]=-E_{w}^{(o)}\cdot h_{x}+\log[\sum_{i\in\mathbb{V}}e^{E_{i}^{(o)}\cdot h_{x}}].$
When we calculate the averaged output token probability $\alpha_{w,\mathcal{D},\theta}=\mathbb{E}_{x\in\mathcal{D}}[P_{\theta}(w|x)]$ of token $w$ on a dataset $\mathcal{D}$ (detecting dataset):
$\begin{split}&-\log\alpha_{w,\mathcal{D},\theta}\approx\mathbb{E}_{x\in\mathcal{D}}[-\log[P_{\theta}(w|x)]]\\ =&-\mathbb{E}_{x\in\mathcal{D}}[E_{w}^{(o)}\cdot h_{x}]+\mathbb{E}_{x\in\mathcal{D}}[\log(\sum_{i\in\mathbb{V}}e^{E_{i}^{(o)}\cdot h_{x}})]\\ =&-E_{w}^{(o)}\cdot\mathbb{E}_{x\in\mathcal{D}}[h_{x}^{(o)}]+\mathbb{E}_{x\in\mathcal{D}}[\log(\sum_{i\in\mathbb{V}}e^{E_{i}^{(o)}\cdot h_{x}})].\end{split}$ (1)
We make a local linear approximation in the approximately equal sign, while we confirm it is precise when the output logits are concentrated in Appendix B.2. Notice that if the LM head is biased, the bias term can be re-constructed equally by fixing one dimension of $h_{x}$ to $1$, w.l.o.g.
We denote $A_{\mathcal{D}}=-\mathbb{E}_{x\in\mathcal{D}}[h_{x}]$, and $B_{\mathcal{D}}=\mathbb{E}_{x\in\mathcal{D}}[\log(\sum_{i\in\mathbb{V}}e^{E_{i}^{(o)}\cdot h_{x}})]$. Here we do another approximation that the $B_{\mathcal{D}}$ is independent to $E_{w}^{(o)}$. This approximation is precise when the logits of the output token $w$ are small, and the vocabulary size $|\mathbb{V}|$ is large. Then, we get an approximated log-linear form between the $\alpha_{w,\mathcal{D},\theta}$ and the $E_{w}^{(o)}$:
@@
With such a derivation, we find that the phenomenon shown in Fig. 1 is the nature of $\mathrm{softmax}$ output head if it has plenty of classification categories, small and concentrated logits to make the approximations numerically accurate. LMs have a very wide output space and undergo regularized training, which makes LMs meet the requirements to have an accurate log-linear correlation between output probabilities and output embeddings. That is, the output probabilities are encoded within a common direction of output embedding.
@@ 2.2 Empirical Confirmation
We conduct experiments to prove our derivation in Eq. 2 empirically correct. First, following Kobayashi et al. (2023), we calculate the averaged output probabilities on an 8192-length sample of shuffled WikiDpr dataset Karpukhin et al. (2020). In detail, we input the sampled data points into the model, and average the output probability distribution among every time step of every input sequence as the token probability distribution of the dataset (see Appendix B.1). Then, we conduct MLR to fit the $A_{\mathcal{D}}$ and the $B_{\mathcal{D}}$.
We run such experiment on GPT2, GPT2-XL Radford et al. (2019), Pythia 2.8B Biderman et al. (2023), and GPT-J Wang and Komatsuzaki (2021). The fitting results are shown in Fig. 2, where good fittings are observed. We also list the adjusted $R^{2}$ i.e. the goodness of fitting in Table 1. Surprisingly, both input and output embeddings have strong correlations with the token probabilities.
More interestingly, as shown in Fig. 3, we find that only the top principal components and a few dimensions in the output embedding are highly correlated to the output token probabilities. In detail, we calculate the absolute Spearman $r$ between the output embedding’s sorted principle components / original dimensions against the output probabilities, and we find that the overwhelming majority of both kinds of dimensions have poor correlations with the output probabilities.
3 Token Probability Editing on Output Embeddings
Our findings suggest that the output probabilities are encoded in a common direction in the output embeddings. So we try to edit the token probabilities by modifying a small portion of dimensions in the output embeddings following such a direction, as an application and further empirical proof of the findings.

### v2

@@ Abstract
In this paper, we investigate the output token probability information in the output embedding of language models. We find an approximate common log-linear encoding of output token probabilities within the output embedding vectors and empirically demonstrate that it is accurate and sparse. As a causality examination, we steer the encoding in output embedding to modify the output probability distribution accurately. Moreover, the sparsity we find in output probability encoding suggests that a large number of dimensions in the output embedding do not contribute to causal language modeling. Therefore, we attempt to delete the output-unrelated dimensions and find more than 30% of the dimensions can be deleted without significant movement in output distribution and sequence generation. Additionally, in the pre-training dynamics of language models, we find that the output embeddings capture the corpus token frequency information in early steps, even before an obvious convergence of parameters starts.
@@
[…]
@@ 1 Introduction
Modern Language Models (LMs) have two kinds of token embeddings: one is the input embedding $E^{(i)}$ located at the earliest layer of LMs, for mapping the input token index into distributed inner representation, the other is the output embedding $E^{(o)}$ in the Language Modeling head (LM head), for mapping the hidden state to the predicted probability distribution of the next token in the causal language modeling task.
@@
Since the output embeddings were often tied with the input embeddings Chung et al. (2020); Press and Wolf (2017), i.e. input embeddings are directly used as the output embedding, the behaviors and features of independent output embeddings are rarely investigated. Along with the scaling of LMs, such embedding tying, which is proven to be harmful to model performance Chung et al. (2020), is gradually being discarded in modern LMs such as GPT-J Wang and Komatsuzaki (2021) and LLaMa 2 Touvron et al. (2023). This raises attention to the independent output embeddings, and explaining its underlying mechanism can be beneficial to understanding and improving LMs.
@@
Figure 1: The PCA result of the output embedding vectors of GPT2. Colors refer to the ranking percentile of the averaged output token probability.
@@
The most obvious and expected role of the LM heads, where the output embedding is located, is to map the last hidden states into token probabilities. Therefore, following Kobayashi et al. (2023), who found an encoding of the averaged output token probability distribution in the bias term of the output LM head, this paper also focuses on the averaged output probabilities [^1] We may omit the “averaged” in the following. as an overall representation of LM outputs. We observe there is a linear-like correlation between the output token probabilities and the output embedding as shown in Fig. 1. In §2, our mathematical derivation indicates that $\mathrm{softmax}$-based LM head naturally encodes the output probabilities log-linearly in a common direction of the output embeddings, as long as the output dimension is sufficiently large, and the output values are concentrated.
@@
To prove our derivation empirically, we conduct Multiple Linear Regression (MLR) on output probabilities against output embeddings, where a strong log-linear correlation is observed, as shown in Fig. 3. Additionally, we find that: (1) Almost all directions highly correlated with output probability are the top principal components of the embedding matrix; (2) Only a few dimensions of output embedding are highly correlated to output probability.
@@
Based on such findings, to further examine the causality between the encoding and the output probabilities, in §3, we try to steer the output probability by the output embedding. We use a linear and sparse vectorized algorithm, modifying a small portion of dimensions along the detected encoding direction in the output embeddings to steer the probabilities. Our experiments find that: even on embedding-tied models, our steering method has respectable precision and a large available range, stably scaling the probability of tokens up to 20x (both scaling up or down), with little disturbing on the normal inference process of LMs. Moreover, with the encoding direction estimated from few-shot examples, the steering remains precise. Such results suggest that the log-linear correlation has good stability and generalization.
@@
Moreover, the sparsity of the probability encoding demonstrates that most of the dimensions of the output embedding have minor effects on output. Therefore, we try removing these dimensions to prune the parameters in output embedding without obvious harm to the causal language modeling in §4. Our experiments show that more than 30% of the output embedding dimensions can be removed without significant impairment.
@@
Our contributions can be summarized as:
  • We find a log-linear correlation as an encoding of the output probability in a particular common and sparse direction of the output embeddings.
  • As a causality examination of the findings, we steer the averaged output probabilities along the detected encoding direction, and remove dimensions with weak correlation to output probabilities without harm to the LMs.
  • Based on the findings, we find that the LMs catch the token frequency in the training corpus at very early training steps.
@@
Model Decoder-only Encoder-only Encoder-decoder
GPT2 GPT2-XL Pythia GPT-J BERT-base BERT-large BART-base BART-large
Parameter # 137M 1.6B 2.8B 6B 110M 340M 139M 406M
Embedding tied
Random Adj.$R^{2}$ $0.001$ $0.000$ $0.001$ $0.000$- -- -- --
Adj.$R^{2}$ on $E^{(o)}_{w}$ $0.892$ $0.893$ $0.856$ $0.882$ $0.833$ $0.826$ $0.998$ $0.952$
Adj.$R^{2}$ on $E^{(i)}_{w}$ $(0.892)$ $(0.893)$ $0.814$ $0.658$- -- -- --
(a) GPT2, 137M, Principal
(b) GPT2-XL, 1.6B, Principal
(c) Pythia, 2.8B, Principal
(d) GPT-J, 6B, Principal
@@
Figure 2: Only a few directions/dimensions of output embedding are strongly correlated to the output probabilities. (a-d): horizontal axis: the principle components of output embedding, vertical axis: absolute Spearman $r$ between the principle and the output probability distribution, color bar: the variance ratio loaded in the principal component; (e-h): horizontal axis: original dimensions, vertical axis: absolute MLR slopes between this dimension and the output probability distribution, color bar: the absolute Spearman correlations on the dimension.
@@ 2 Token Probability Encoding in Output Embedding
In this section, we preliminarily reveal how the output probabilities are encoded in the output embedding by mathematical derivation, then conduct numerical experiments to confirm it empirically.
Input: Language model $\mathrm{P}_{\theta}(x)$ with output embedding $E^{(o)}$; Probability detecting dataset $\mathcal{D}$;
Token index to be steered $w$; Expected scale of the steered token’s probability $r$;
Steering amount allocation parameter $b$
@@
$\mathbf{\alpha}_{\mathcal{D},\theta}\leftarrow|\mathcal{D}|^{-1}\sum_{x\in\mathcal{D}}\mathrm{P}_{\theta}(x)$ ;
  /* Calculate the averaged probability distribution */
@@
$A_{\mathcal{D}},\mathcal{S}\leftarrow\mathrm{MLR}\left(-\log\mathbf{\alpha}_{\mathcal{D},\theta},E^{(o)}\right)$ ;
  /* Conduct MLR, get the slope and significance */
@@
$\Omega\leftarrow|A_{\mathcal{D}}|^{b}\times\mathcal{S}\times\left\||A_{\mathcal{D}}|^{b}\times\mathcal{S}\right\|_{1}^{-1}$ ;
  /* Allocate the steering weight to elements */
@@
  /* Apply the steering to output embedding */
@@
10return Updated LM $\mathrm{P}_{\theta^{\prime}}(x)$ with new output embedding $E_{w}^{{}^{\prime}(o)}$
2 $\mathbf{\alpha}_{\mathcal{D},\theta}\in\mathbb{R}^{|\mathbb{V}|}$: the output token probability distribution among the vocabulary; $|\mathcal{D}|$: the length of the probability detecting dataset; $\mathcal{S}$: the element-wise correlation significance ($p$-value) of MLR;
@@
Algorithm 1 Output token probability steering.
4 $\mathbf{\alpha}_{\mathcal{D},\theta}\in\mathbb{R}^{|\mathbb{V}|}$: the output token probability distribution among the vocabulary; $|\mathcal{D}|$: the length of the probability detecting dataset; $\mathcal{S}$: the element-wise correlation significance ($p$-value) of MLR;
@@ 2.1 Mathematical Log-linear Form
$\mathrm{P}_{\theta}(w|x)=\frac{e^{E_{w}^{(o)}\cdot h_{x}}}{\sum_{i\in\mathbb{V}}e^{E_{i}^{(o)}\cdot h_{x}}},$
@@
$-\log\left[\mathrm{P}_{\theta}(w|x)\right]=-E_{w}^{(o)}\cdot h_{x}+\log\left[\sum_{i\in\mathbb{V}}e^{E_{i}^{(o)}\cdot h_{x}}\right].$
When we calculate the averaged output token probability $\alpha_{w,\mathcal{D},\theta}=\mathbb{E}_{x\in\mathcal{D}}[\mathrm{P}_{\theta}(w|x)]$ of token $w$ on a dataset $\mathcal{D}$ (detecting dataset):
$\begin{split}&-\log\alpha_{w,\mathcal{D},\theta}\approx\mathbb{E}_{x\in\mathcal{D}}[-\log[\mathrm{P}_{\theta}(w|x)]]\\ =&-\mathbb{E}_{x\in\mathcal{D}}[E_{w}^{(o)}\cdot h_{x}]+\mathbb{E}_{x\in\mathcal{D}}[\log(\sum_{i\in\mathbb{V}}e^{E_{i}^{(o)}\cdot h_{x}})]\\ =&-E_{w}^{(o)}\cdot\mathbb{E}_{x\in\mathcal{D}}[h_{x}^{(o)}]+\mathbb{E}_{x\in\mathcal{D}}[\log(\sum_{i\in\mathbb{V}}e^{E_{i}^{(o)}\cdot h_{x}})].\end{split}$ (1)
We make a local linear approximation in the approximately equal sign, while we confirm it is precise when the output logits are concentrated in Appendix A.2. Notice that if the LM head is biased, the bias term can be re-constructed equally by fixing one dimension of $h_{x}$ to $1$, w.l.o.g.
We denote $A_{\mathcal{D}}=-\mathbb{E}_{x\in\mathcal{D}}[h_{x}]$, and $B_{\mathcal{D}}=\mathbb{E}_{x\in\mathcal{D}}\left[\log\left(\sum_{i\in\mathbb{V}}e^{E_{i}^{(o)}\cdot h_{x}}\right)\right]$. Here we make another approximation that the $B_{\mathcal{D}}$ is independent to $E_{w}^{(o)}$, which is precise when the logits of the output token $w$ are small, and the vocabulary size $|\mathbb{V}|$ is large. Then, we get an approximated log-linear form between the $\alpha_{w,\mathcal{D},\theta}$ and the $E_{w}^{(o)}$:
@@
With such a derivation, we find that the phenomenon shown in Fig. 1 is the nature of $\mathrm{softmax}$-based output head if it has plenty of output dimensions, and small and concentrated logits to make the approximations numerically accurate. LMs have a very wide output space and undergo regularized training, which makes LMs meet the requirements to have an accurate log-linear correlation between output probabilities and output embeddings. That is, the output probabilities are encoded within a common direction of output embedding.
@@ 2.2 Empirical Confirmation
(a) GPT2, 137M
(b) GPT-J, 6B
Figure 3: The MLR results on GPT2 and GPT-J.
We conduct experiments to prove our derivation in Eq. 2 empirically accurate. First, following Kobayashi et al. (2023), we define the detect dataset $\mathcal{D}$ as an 8192-sample of shuffled WikiDpr dataset Karpukhin et al. (2020) to calculate the averaged output probabilities. In detail, we input the sampled data points into the model, and average the output probability distribution among every time step (aligned with input token) of every input sequence as the token probability distribution of the dataset (see Appendix A.1). Then, we conduct MLR to fit the $A_{\mathcal{D}}$ and the $B_{\mathcal{D}}$.
We conduct such experiment on GPT2, GPT2-XL Radford et al. (2019), Pythia 2.8B Biderman et al. (2023), and GPT-J Wang and Komatsuzaki (2021). The fitting results are shown in Fig. 3, where good fittings are observed. We also list the adjusted $R^{2}$ i.e. the goodness of fitting in Table 1. Surprisingly, both input and output embeddings have strong correlations with the token probabilities. Additionally, as a reference, we provide some rough results for the encoder-only (BERT-base and BERT-large Kenton and Toutanova (2019)) and encoder-decoder (BART-base and BART-large Lewis (2019)) models, in addition to the decoder-only model [^2] Note: This paper focuses exclusively on an in-depth discussion of the decoder-only model..
More interestingly, as shown in Fig. 2, we find that only the top principal components and a few dimensions in the output embedding are highly correlated to the output token probabilities. That is, such probability encoding is sparse. In detail, we calculate the correlation coefficient (absolute Spearman $r$ / MLR slope) between the output embedding’s sorted principle components / original dimensions against the output probabilities, and we find that the overwhelming majority of both kinds of dimensions have poor correlations with the output probabilities.
3 Token Probability Steering on Output Embeddings
Our findings in §2 suggest that the output probabilities are encoded in a common direction in the output embeddings. So, as a causality examination, we steer the output embeddings along this detected direction and confirm that (1) such steering can control the output probabilities accurately on a large scale, implying a robust log-linear token probability encoding. (2) steering based on a few-shot detect dataset $\mathcal{D}$ can still be effective, confirming that such encoding is significant and widely present.